A person works and earns $3360 dollars but on a pay stub pays 24% to tax. How much money is left over after taxes? Round your answer to the nearest cent.
Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52781)

.

Left over after taxes = 3360 - 0.24*3360  dollars.


You do the calculations.


Answer by Theo(13342)

3360 minus .24 * 3360 = 3360 * (1 - .24) = 3360 * .76 = 2553.6.
2553.6 is what is left over after taxes have been withheld.